Two workers can trim 400 meters of hedge in 10 hours. how long will it take 5 workers to trim 600 meters of hedge?
______hours.
Thank You.

Answer by stanbon(75887)

Two workers can trim 400 meters of hedge in 10 hours. how long will it take 5 workers to trim 600 meters of hedge?
2 workers = 400meters/ 10 hrs.
Multiply thru by (5/2) to get:
5 workers = 1000 meters/ 10 hrs.
600/1000= (3/5)= 0.6
5 workers = [0.6(1000/meters)/0.6(10 hr.)]
5 workers = 600 meters/ 6 hrs.
Cheers,
Stan H.